### v2

# Finding All Solutions of qKZ Equations in Characteristic $p$

[…]
Finding All Solutions of qKZ Equations in Characteristic $p$
[…]
Finding All Solutions of qKZ Equations
in Characteristic $\boldsymbol{p}$
[…]
Evgeny MUKHIN a and Alexander VARCHENKO b
[…]
E. Mukhin and A. Varchenko
[…]
a) Department of Mathematical Sciences, Indiana University Indianapolis,
[…]
b) Department of Mathematics, University of North Carolina at Chapel Hill,
[…]
Received September 23, 2025, in final form December 16, 2025; Published online January 02, 2026
[…]
In [J. Lond. Math. Soc. 109 (2024), e12884, 22 pages], the difference qKZ equations were considered modulo a prime number $p$ and a family of polynomial solutions of the qKZ equations modulo $p$ was constructed by an elementary procedure as suitable $p$-approximations of the hypergeometric integrals. In this paper, we study in detail the first family of nontrivial examples of the qKZ equations in characteristic $p$. We describe all solutions of these qKZ equations in characteristic $p$ by demonstrating that they all stem from the $p$-hypergeometric solutions. We also prove a Lagrangian property (called the orthogonality property) of the subbundle of the qKZ bundle spanned by the $p$-hypergeometric sections. This paper extends the results of [arXiv:2405.05159] on the differential KZ equations to the difference qKZ equations.
[…]
qKZ equations; $p$-hypergeometric solutions; orthogonality relations; $p$-curvature
[…]
11D79; 12H25; 32G34; 33C05; 33E30
1 Introduction
The Knizhnik–Zamolodchikov (KZ) differential equations are a system of linear differential equations, satisfied by conformal blocks on the sphere in the WZW model of conformal field theory, see [4]. The quantum Knizhnik–Zamolodchikov (qKZ) equations are a difference version of the KZ equations which naturally appear in the representation theory of Yangians (rational case) and quantum affine algebras (trigonometric case), see [1, 3]. The qKZ equations may be regarded as a deformation of the KZ differential equations.
As a rule one considers the KZ and qKZ equations over the field of complex numbers. Then these differential and difference equations are solved in multidimensional hypergeometric integrals.
In [6], the differential KZ equations were considered modulo a prime integer $p$. It turned out that modulo $p$ the KZ equations have a family of polynomial solutions. The construction of these solutions was analogous to the construction of the multidimensional hypergeometric solutions, and these polynomial solutions were called the $p$-hypergeometric solutions.
In [5], the rational $\mathfrak{sl}_{2}$} qKZ equations with values in the $n$-th tensor power of the vector representation $L$ and an integer step $\kappa$ were considered modulo $p$. A family of polynomial solutions modulo $p$ of these equations was constructed and called the $p$-hypergeometric solutions.
In this paper, we address the problem of whether all solutions of the qKZ equations in characteristic $p$ are generated by the $p$-hypergeometric solutions. We consider the first family of nontrivial examples of the qKZ equations and demonstrate that, indeed, in this case, all solutions of the qKZ equations stem from the $p$-hypergeometric solutions.
Let $\mathbb{K}$} be a field of characteristic $p$. The qKZ equations for a function $f(z_{1},\dots,z_{n})$ with values in the $\mathbb{K}$}-vector space $L^{\otimes n}$ and step $\kappa\in{\mathbb{K}}^{\times}$ have the form
$f(z_{1},\dots,z_{a}-\kappa,\dots,z_{n})=K_{a}(z;\kappa)f(z),\qquad a=1,\dots,n,$
where the linear operators $K_{a}(z;\kappa)$ are given in terms of the rational $\mathfrak{sl}_{2}$} $R$-matrix, see (2.2). The operators $K_{a}(z;\kappa)$ commute with the diagonal action of $\mathfrak{sl}_{2}$}, and, therefore, it is sufficient to solve the qKZ equations only with values in the space of singular vectors of a given weight. In this paper, we study the qKZ equations with values in $V:=\operatorname{Sing}L^{\otimes n}[n-2]\subset L^{\otimes n}$, the subspace of singular vectors of weight n2n-2. We have $\dim V=n-1$.
There are two cases: $\kappa\in{\mathbb{K}}\setminus{\mathbb{F}}_{p}$ and $\kappa\in{\mathbb{F}}_{p}^{\times}$.
Theorem 1.1.
Let $p$ be a prime number that does not divide $n$. For $\kappa\in{\mathbb{K}}\setminus{\mathbb{F}}_{p}$, there does not exist a nonzero rational $V$-valued function $f(z_{1},\dots,z_{n})$ which is a solution of the qKZ equations with parameter $\kappa$.
See Corollary 7.12.
Assume that $\kappa\in{\mathbb{F}}_{p}^{\times}$. Let $0<k<p$ be the positive integer such that $\kappa k\equiv-1\pmod{p}$}. Let $[x]$ denote the integer part of a real number and $d(\kappa):=\big[\frac{kn}{p}\big]$. If $p$ does not divide $n$, then $d(\kappa)+d(-\kappa)=n-1=\dim V$}.
In [5], we constructed $d(\kappa)$ $V$-valued $p$-hypergeometric solutions of the qKZ equations denoted $Q^{\ell p-1}(z;\kappa)$}, $\ell=1,\dots,d(\kappa)$. In this paper, we show that these solutions are linearly independent over the field ${\mathbb{K}}(z_{1},\dots,z_{n})$, see Theorem 5.6.
Theorem 1.2.
Let $p>n$, $\kappa\in{\mathbb{F}}_{p}^{\times}$, and $0<d(\kappa)<n-1$. Let $f(z)$ be a $V$-valued rational function in $z$ which is a solution the qKZ equations with step $\kappa$. Then $f(z)$ is a linear combination of the $p$-hypergeometric solutions $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$, with coefficients which are scalar rational functions in $z_{i}^{p}-z_{i}$, $i=1,\dots,n$.
See Theorem 7.9. Notice that $h(x)=x^{p}-x\in{\mathbb{K}}[x]$ is a 1-periodic polynomial, $h(x+1)=h(x)$. In particular, $h(x+\kappa)=h(x)$.
If $d(\kappa)=n-1$ or 0, all solutions of the qKZ equations with values in $V$ and step $\kappa$ are described in Section 7.4.
We prove the orthogonality relations for $p$-hypergeometric solutions of the qKZ equations with steps $\kappa$ and $\kappa$.
Theorem 1.3.
Let $p>n$ and $0<d(\kappa)<n-1$. Then for any $\ell\in\{1,\dots,d(\kappa)\}$ and $m\in\{1,\dots,d(-\kappa)\}$, we have
$S\big(Q^{mp-1}(-z;-\kappa),Q^{\ell p-1}(z;\kappa)\big)=0,$
where $S$ is the Shapovalov form.
See Theorem 6.1.
Define the $p$-curvature operators of the qKZ equations with values in $V$ and step $\kappa$ by the formula
$C_{a}(z_{1},\dots,z_{n};\kappa):=K_{a}(z_{1},\dots,z_{a}-(p-1)\kappa,\dots,z_{n};\kappa)$
[…]
for $a=1,\dots,n$, and the reduced $p$-curvature operators by the formula
$\hat{C}_{a}(z_{1},\dots,z_{n};\kappa):=C_{a}(z_{1},\dots,z_{n};\kappa)-1.$
If $f(z_{1},\dots,z_{n})$ is a solution of the qKZ equations, then $\hat{C}_{a}f=0$ for all $a$.
Theorem 1.4.
Let $p>n$. Then the reduced $p$-curvature operators have the following properties:
  1. $\operatorname{(i)}$
    If $\kappa\in{\mathbb{K}}\setminus{\mathbb{F}}_{p}$, then all reduced $p$-curvature operators $\hat{C}_{a}(z;\kappa)$, $a=1,\dots,n$, are nondegenerate for generic $z$.
  2. $\operatorname{(ii)}$
    If $\kappa\in{\mathbb{F}}_{p}^{\times}$ and $d(\kappa)=n-1$ or 0, then all reduced $p$-curvature operators $\hat{C}_{a}(z;\kappa)$, $a=1,\dots,n$, are equal to zero.
  3. $\operatorname{(iii)}$
    If $\kappa\in{\mathbb{F}}_{p}^{\times}$ and $0<d(\kappa)<n-1$, then all reduced $p$-curvature operators $\hat{C}_{a}(z;\kappa)$, $a=1,\dots,n$, are nonzero. For every $a$, the span of the $p$-hypergeometric solutions $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$, lies in the kernel of $\hat{C}_{a}(z,\kappa)$ and contains the image of $\hat{C}_{a}(z;\kappa)$. Also, for all $a$, $b$,
    $\hat{C}_{a}(z;\kappa)\hat{C}_{b}(z;\kappa)=0,\qquad\hat{C}_{a}(z,-\kappa)+\hat{C}_{a}(-z;\kappa)^{*}=0,$
    where for an operator $T\colon V\to V$ we denote by $T^{*}$ the operator dual to $T$ under the Shapovalov form.
See Theorem 7.3 and Lemmas 7.7, 7.11.
In a suitable limit the difference qKZ equations on $L^{\otimes n}$ degenerate to the differential KZ equations on $L^{\otimes n}$,
$\kappa\frac{\partial f}{\partial z_{a}}=\sum_{j\neq a}\frac{P^{(a,j)}-1}{z_{a}-z_{j}}f,\qquad a=1,\dots,n,$
where $P^{(a,j)}$ is the permutation operator of the $a$-th and $j$-th tensor factors of $L^{\otimes n}$.
In [8], the differential KZ equations over a field $\mathbb{K}$} of characteristic $p$ with values in $V\subset L^{\otimes n}$ were studied in detail. Our paper extends the results of [8] from the differential KZ equations to the difference qKZ equations. The proofs of Theorems 1.1, 1.2, and 1.4 are based on the corresponding results in [8] for the differential KZ equations.
On the differential and difference equations in characteristic $p$ and associated $p$-curvature see also [2, 9].
2 Difference qKZ equations
2.1 Notations
In this paper, $p$ is a prime and $\mathbb{K}$} a field of characteristic $p$.
Consider the Lie algebra $\mathfrak{sl}_{2}$} over $\mathbb{K}$} with basis $e$, $f$, $h$ and relations $[e,f]=h$, $[h,e]=2e$, [h,f]=2f[h,f]=-2f. Let $L$ be the two-dimensional $\mathfrak{sl}_{2}$}-module with basis $v_{1}$, $v_{2}$ and the action $ev_{1}=0$, $ev_{2}=v_{1}$, $fv_{1}=v_{2}$, $fv_{2}=0$, $hv_{1}=v_{1}$, $hv_{2}=-v_{2}$.
For a positive integer $n>1$, consider the $\mathfrak{sl}_{2}$}-module $L^{\otimes n}$.
Let $\mathcal{I}_{l}$ be the set of all $l$-element subsets of $\{1,\dots,n\}$. For a subset $I\subset\{1,\dots,n\}$, denote
$v_{I}=v_{i_{1}}\otimes\dots\otimes v_{i_{n}}\in L^{\otimes n},$
where $i_{j}=2$ if $i_{j}\in I$ and $i_{j}=1$ if $i_{j}\notin I$. Denote by $L^{\otimes n}[n-2l]$ the span of the vectors $\{v_{I}\mid I\in\mathcal{I}_{l}\}$. We have a direct sum decomposition,
$L^{\otimes n}=\bigoplus_{l=0}^{n}L^{\otimes n}[n-2l].$
Let $\operatorname{Sing}L^{\otimes n}[n-2l]\subset L^{\otimes n}[n-2l]$ be the subspace of singular vectors (the vectors annihilated by $e$).
2.2 qKZ equations
Define the rational $R$-matrix acting on $L^{\otimes 2}$, $R(u)=\frac{u-P}{u-1}$, where $P$ is the permutation of tensor factors of $L^{\otimes 2}$. The $R$-matrix satisfies the Yang–Baxter and unitarity equations,
$R^{(12)}(u-v)R^{(13)}(u)R^{(23)}(v)=R^{(23)}(v)R^{(13)}(u)R^{(12)}(u-v),$
$R^{(12)}(u)R^{(21)}(-u)=1.$ (2.1)
The first equation is an equation in $\operatorname{End}\big(L^{\otimes 3}\big)$. The superscript indicates the factors of $L^{\otimes 3}$ on which the corresponding operators act.
Let $z=(z_{1},\allowbreak\dots,\allowbreak z_{n})$. Define the qKZ operators $K_{1},\dots,K_{n}\mskip 2.00002mu$ acting on $L^{\otimes n}$
$K_{a}(z;\kappa)=R^{(a,a-1)}(z_{a}-z_{a-1}-\kappa)\cdots R^{(a,1)}(z_{a}-z_{1}-\kappa)$
[…]
where $\kappa\in{\mathbb{K}}^{\times}$ is a parameter.
The qKZ operators preserve the weight decomposition of $L^{\otimes n}$, commute with the $\mathfrak{sl}_{2}$}-action, and form a discrete flat connection with step $\kappa$ on the trivial bundle $L^{\otimes n}\times{\mathbb{K}}^{n}\to{\mathbb{K}}^{n}$,
$K_{a}(z_{1},\dots,z_{b}-\kappa,\dots,z_{n};\kappa)K_{b}(z;\kappa)=K_{b}(z_{1},\dots,z_{a}-\kappa,\dots,z_{n};\kappa)K_{a}(z;\kappa)$
for $a,b=1,\dots,n$, see [3].
The system of difference equations with step $\kappa$,
$s(z_{1},\dots,z_{a}-\kappa,\dots,z_{n})=K_{a}(z;\kappa)s(z),\qquad a=1,\dots,n,$ (2.3)
for an $L^{\otimes n}$-valued function $s(z)$, is called the qKZ equations with step $\kappa$.
Since the qKZ operators commute with the action of $\mathfrak{sl}_{2}$} on $L^{\otimes n}$, the qKZ operators preserve the subbundle $\operatorname{Sing}L^{\otimes n}[n-2l]\times{\mathbb{K}}^{n}\to{\mathbb{K}}^{n}$ for every integer $l$.
Define the translation operators $\bf T_{a}$ by
$({\bf T}_{a}f)(z_{1},\dots,z_{a},\dots,z_{n})=f(z_{1},\dots,z_{a}-\kappa,\dots,z_{n}).$
The difference operators $\nabla_{a}={\bf T}_{a}^{-1}K_{a}(z;\kappa)$ are called the connection operators of the qKZ difference connection. We have $\nabla_{a},\nabla_{b}]=0$.
2.3 $\boldsymbol{p}$-curvature of the qKZ connection
Define the $p$-curvature operators of the qKZ connection by
$C_{a}(z;\kappa)=K_{a}(z_{1},\dots,z_{a}-(p-1)\kappa,\dots,z_{n};\kappa)K_{a}(z_{1},\dots,z_{a}-(p-2)\kappa,\dots,z_{n};\kappa)\cdots$
[…]
for $a=1,\dots,n$. In other words, $C_{a}=(\nabla_{a})^{p}$.
For every $a$, the operator $C_{a}(z;\kappa)$ acts on fibers of the bundle $L^{\otimes n}\times{\mathbb{K}}^{n}\to{\mathbb{K}}^{n}$ and defines an endomorphism of the qKZ connection,
$K_{b}(z_{1},\dots,z_{n};\kappa)C_{a}(z_{1},\dots,z_{n};\kappa)=C_{a}(z_{1},\dots,z_{b}-\kappa,\dots,z_{n};\kappa)K_{b}(z_{1},\dots,z_{n};\kappa).$
The operators $C_{a}(z;\kappa)$ commute, i.e., $C_{a}(z;\kappa),C_{b}(z;\kappa)]=0$.
If $s(z;\kappa)$ is a flat section of the qKZ discrete connection,
$s(z_{1},\dots,z_{a}-\kappa,\dots,z_{n};\kappa)=K_{a}(z;\kappa)s(z;\kappa),\qquad a=1,\dots,n,$
then $s(z;\kappa)$ is an eigenvector of the $p$-curvature operators with eigenvalue 1,
$s(z;\kappa)=C_{a}(z;\kappa)s(z;\kappa),\qquad a=1,\dots,n.$ (2.4)
An operator $C_{a}(z;\kappa)$ is a rational function in $z$ with the denominator
$D_{a}(z;\kappa)={}$ $\prod_{j\neq a}\prod_{m=0}^{p-1}(z_{a}-z_{j}-m\kappa-1)$
$={}$ $\prod_{j\neq a}\big(z_{a}^{p}-\kappa^{p-1}z_{a}+(-z_{j})^{p}+\kappa^{p-1}z_{j}+(-1)^{p}+\kappa^{p-1}\big).$
It is convenient to introduce the reduced $p$-curvature operators by the formula
$\hat{C}_{a}(z;\kappa)=C_{a}(z;\kappa)-1,$ (2.5)
and the normalized $p$-curvature operators by the formula
$\tilde{C}_{a}(z;\kappa)=D_{a}(z;\kappa)(C_{a}(z;\kappa)-1).$
The normalized $p$-curvature operators are polynomials in $z$ of degree $\leqslant(n-1)p$.
2.4 Differential KZ equations
For $\kappa\in{\mathbb{K}}^{\times}$, the differential KZ operators
$\nabla_{a}^{\operatorname{KZ}}=\kappa\frac{\partial}{\partial z_{a}}-\sum_{j\neq a}\frac{P^{(a,j)}-1}{z_{a}-z_{j}},\qquad a=1,\dots,n,$
define a flat KZ connection on $L^{\otimes n}\times K^{n}\to{\mathbb{K}}^{n}$, $\big[\nabla_{a}^{\operatorname{KZ}},\nabla_{b}^{\operatorname{KZ}}\big]=0$. The operators $\nabla_{a}^{\operatorname{KZ}}$ commute with the $\mathfrak{sl}_{2}$}-action on $L^{\otimes n}$. The system of equations
$\kappa\frac{\partial f}{\partial z_{a}}=\sum_{j\neq a}\frac{P^{(a,j)}-1}{z_{a}-z_{j}}f,\qquad a=1,\dots,n,$ (2.6)
is called the differential KZ equations with parameter $\kappa$. The KZ operators preserve every subbundle $\operatorname{Sing}L^{\otimes n}[n-2l]\times{\mathbb{K}}^{n}\to{\mathbb{K}}^{n}$. Denote
$H_{a}(z)=\sum_{j\neq a}\frac{P^{(a,j)}-1}{z_{a}-z_{j}},$
the Gaudin Hamiltonians.
The $p$-curvature operators of the KZ connection are defined by the formula
$C_{a}^{\operatorname{KZ}}(z;\kappa):=\left(\nabla_{a}^{\operatorname{KZ}}\right)^{p}.$
They define an endomorphism of the KZ connection, $\big[C_{a}^{\operatorname{KZ}},\nabla_{b}^{\operatorname{KZ}}\big]=0$.
An operator $C^{\operatorname{KZ}}_{a}(z;\kappa)$ is a rational function in $z$ with the denominator
$D^{\operatorname{KZ}}_{a}(z;\kappa)=\prod_{j\neq a}\big(z_{a}^{p}-z_{j}^{p}\big).$
It is convenient to introduce the normalized $p$-curvature operator by the formula
$\tilde{C}^{\operatorname{KZ}}_{a}(z;\kappa)=D^{\operatorname{KZ}}_{a}(z;\kappa)C_{a}^{\operatorname{KZ}}(z;\kappa).$
The normalized $p$-curvature operator is a homogeneous polynomial in $z$ of degree (n2)p(n-2)p if nonzero.
2.5 KZ equations as a limit of qKZ equations
Let $f(z_{1},\dots,z_{n})$ satisfy the qKZ equations,
$f(z_{1},\dots,z_{a}-\kappa,\dots,z_{n})=K_{a}(z;\kappa)f(z),\qquad a=1,\dots,n.$
Let $\alpha$ be a formal parameter. Define $g(w_{1},\dots,w_{n};\alpha):=f(w_{1}/\alpha,\dots,w_{n}/\alpha)$. Then
$g(w_{1},\dots,w_{a}-\alpha\kappa,\dots,w_{n};\alpha)=K_{a}(w/\alpha;\kappa)g(w;\alpha),\qquad a=1,\dots,n,$ (2.7)
where
$K_{a}(w/\alpha;\kappa)=R^{(a,a-1)}(w_{a}-w_{a-1}-\alpha\kappa;\alpha)\cdots R^{(a,1)}(w_{a}-w_{1}-\alpha\kappa;\alpha)$
[…]
$R(u;\alpha)=\frac{u-\alpha P}{u-\alpha}=1-\alpha\frac{P-1}{u-\alpha}.$
Equation (2.2) gives
$g-\alpha\kappa\frac{\partial g}{\partial w_{a}}+\mathcal{O}\big(\alpha^{2}\big)=\big(1-\alpha H_{a}(w)+\mathcal{O}\big(\alpha^{2}\big)\big)g.$
In the limit $\alpha\to 0$, we obtain the KZ differential equations
$\kappa\frac{\partial g}{\partial w_{a}}(w)=H_{a}(w)g(w).$
Lemma 2.1.
Let $\hat{C}_{a}(z;\kappa)$ be a reduced $p$-curvature operators of the qKZ connection. Then
$C_{a}^{\operatorname{KZ}}(z;\kappa)=\lim_{\alpha\to 0}\hat{C}_{a}(z_{1}/\alpha,\dots,z_{n}/\alpha;\kappa)/\alpha^{p}.$
Proof.
Let $C_{a,\alpha}$ be the $p$-curvature operator of equation (2.7). Clearly,
$C_{a,\alpha}(w;\kappa)=C_{a}(w_{1}/\alpha,\dots,w_{n}/\alpha;\kappa).$
Hence we need to prove that
$C_{a}^{\operatorname{KZ}}(w;\kappa)=\lim_{\alpha\to 0}(C_{a,\alpha}(w;\kappa)-1)/\alpha^{p}.$ (2.8)
Let $\nabla_{a,\alpha}$ be the connection operator of the discrete connection (2.7),
$(\nabla_{a,\alpha}g)(w)=K_{a}(w_{1}/\alpha,\dots,(w_{a}+\alpha\kappa)/\alpha,\dots,w_{n}/\alpha;\kappa)g(w_{1},\dots,w_{a}+\alpha\kappa,\dots,w_{n}).$
Then
$(\nabla_{a,\alpha}g)(w)=g-\alpha H_{a}(w)g+\alpha\kappa\frac{\partial g}{\partial w_{a}}+\mathcal{O}\big(\alpha^{2}\big).$
Hence
$(\nabla_{a,\alpha}-1)g=\alpha\nabla^{\operatorname{KZ}}_{a}g+\mathcal{O}\big(\alpha^{2}\big).$
Thus
$C_{a,\alpha}(w;\kappa)-1=(\nabla_{a,\alpha})^{p}-1=(\nabla_{a,\alpha}-1)^{p}=\alpha^{p}\big(\nabla^{\operatorname{KZ}}_{a}\big)^{p}+\mathcal{O}\big(\alpha^{p+1}\big).$
Corollary 2.2.
The degree of the reduced $p$-curvature operator $\tilde{C}_{a}(z;\kappa)$ as a polynomial in $z$ is not greater than (n2)p(n-2)p. Moreover, we have $\tilde{C}_{a}(z;\kappa)=\tilde{C}_{a}^{\operatorname{KZ}}(z;\kappa)+f_{1}(z)$, where $f_{1}(z)$ is a polynomial in $z$ of degree less than (n2)p(n-2)p.
Proof.
The first statement of the corollary follows from the fact that the denominator of the reduced $p$-curvature operator $\hat{C}_{a}(z;\kappa)$ is of degree (n1)p(n-1)p. Hence the numerator is of degree not greater than (n2)p(n-2)p for the limit in (2.8) to exist. The second statement is clear. ∎
2.6 Limit of solutions
Let $f(z_{1},\dots,z_{n})$ be a polynomial solution of the qKZ equations (2.3). Let $\deg f(z)=d$ and $f(z)=f_{0}(z)+f_{1}(z)$}, where $f_{0}(z)$ is a homogeneous polynomial of degree $d$ and $f_{1}(z)$ is a polynomial of degree less than $d$.
Lemma 2.3.
The polynomial $f_{0}(z)$ is a solution of the KZ equations (2.6).
Proof.
Let $\alpha$ be a formal parameter. Define $g(w_{1},\dots,w_{n};\alpha)=\alpha^{d}f(w_{1}/\alpha,\dots,w_{n}/\alpha)$. Then $g(w_{1},\dots,w_{n};\alpha)=f_{0}(w)+\alpha g_{1}(w;\alpha)$, where $g_{1}(w;\alpha)=\alpha^{d-1}f_{1}(w/\alpha)$ is a polynomial in $w$ and $\alpha$. We have
$g(w_{1},\dots,w_{a}-\alpha\kappa,\dots,w_{n};\alpha)=K_{a}(w/\alpha;\kappa)g(w;\alpha),\qquad a=1,\dots,n.$
Let
$d_{a}(w;\alpha)=\prod_{j=1}^{a-1}(w_{a}-w_{j}-\alpha\kappa-\alpha)\prod_{j=a+1}^{n}(w_{a}-w_{j}-\alpha)$
be the denominator of $K_{a}(w/\alpha;\kappa)$ and
$n_{a}(w;\alpha)=\big(w_{a}-w_{a-1}-\alpha\kappa-\alpha P^{(a,a-1)}\big)\cdots\big(w_{a}-w_{a+1}-\alpha P^{(a,a+1)}\big)$
the numerator. Then (2.7) can be written as a polynomial equation
$d_{a}(w;\alpha)g(w_{1},\dots,w_{a}-\alpha\kappa,\dots,w_{n};\alpha)=n_{a}(w;\alpha)g(w;\alpha)$ (2.9)
in the variables $w$ and $\alpha$. Equation (2.9) gives an equation in $w$ for every fixed power of $\alpha$ in (2.9). The equation corresponding to the first power of $\alpha$ after division by $d_{a}(w;0)$ becomes
$\kappa\frac{\partial f_{0}}{\partial w_{a}}(w)=\sum_{j\neq a}\frac{P^{(a,j)}-1}{w_{a}-w_{j}}f_{0}(w).$
2.7 Dual qKZ equations
Denote $W=L^{\otimes n}$. Let $W^{*}$ be the dual space of $W$, and $\langle,\rangle\colon W^{*}\otimes W\to{\mathbb{K}}$ the canonical pairing. Let $K_{a}^{*}(z;\kappa)\colon W^{*}\to W^{*}$ be the operators dual to the operators $K_{a}(z;\kappa)$. Denote $\tilde{K}_{a}(z;\kappa)=(K_{a}^{*}(z;\kappa))^{-1}$.
We have
$\tilde{K}_{a}(z_{1},\dots,z_{b}-\kappa,\dots,z_{n};\kappa)\tilde{K}_{b}(z;\kappa)=\tilde{K}_{b}(z_{1},\dots,z_{a}-\kappa,\dots,z_{n};\kappa)\tilde{K}_{a}(z;\kappa)$
for all $a$, $b$, and also
$\langle x,y\rangle=\big\langle\tilde{K}_{a}(z;\kappa)x,K_{a}(z;\kappa)y\big\rangle$ (2.10)
for all $a$ and $x\in W^{*},y\in W$.
The system of difference equations with step $\kappa$,
$\tilde{s}(z_{1},\dots,z_{a}-\kappa,\dots,z_{n})=\tilde{K}_{a}(z;\kappa)\tilde{s}(z),\qquad a=1,\dots,n,$ (2.11)
for an $W^{*}$-valued function $\tilde{s}(z)$, is called the dual qKZ equations.
If $s(z)$ is a solution of the qKZ equations (2.3) and $\tilde{s}(z)$ is a solution of the dual qKZ equations (2.11), then the function $\langle\tilde{s}(z),s(z)\rangle$ is $\kappa$-periodic with respect to every $z_{a}$,
$\langle\tilde{s}(z_{1},\dots,z_{a}-\kappa,\dots,z_{n}),s(z_{1},\dots,z_{a}-\kappa,\dots,z_{n},)\rangle=\langle\tilde{s}(z),s(z)\rangle.$
The set of vectors $\{v_{I}\mid I\subset\{1,\dots,n\}\}$ is a basis of $W$. Define the nondegenerate symmetric bilinear form $S$ on $W$ by the formula $S(v_{I},v_{J})=\delta_{I,J}$. The form is called the (tensor) Shapovalov form. We identify $W^{*}$ and $W$ with the help of the Shapovalov form.
Under this identification, the operators $R^{(i,j)}(u)$ become symmetric. Using (2.1) we obtain the following formula for the operators $\tilde{K}_{a}(z;\kappa)$ as operators on $W$,
$\tilde{K}_{a}(z;\kappa)=R^{(a,a-1)}(-z_{a}+z_{a-1}+\kappa)\cdots R^{(a,1)}(-z_{a}+z_{1}+\kappa)$
[…]
In other words,
$\tilde{K}_{a}(z_{1},\dots,z_{n};\kappa)=K_{a}(-z_{1},\dots,-z_{n};-\kappa),$
$S(K_{a}(-z_{1},\dots,-z_{n};-\kappa)x,K_{a}(z_{1},\dots,z_{n};\kappa)y)=S(x,y)$
for all $x,y\in W$.
Now the dual qKZ equations for a $W$-valued function $\tilde{s}(z)$ take the form
$\tilde{s}(z_{1},\dots,z_{a}-\kappa,\dots,z_{n})=K_{a}(-z_{1},\dots,-z_{n};-\kappa)s(z),\qquad a=1,\dots,n.$
These formulas prove the following lemma.
Lemma 2.4.
Let $s(z)$ be a solution of the qKZ equations (2.3) with step $\kappa$ and $\tilde{s}(z)$ a solution of the qKZ equations (2.3) with step $\kappa$. Then the function $S(\tilde{s}(-z),s(z))$ is $\kappa$-periodic with respect to every variable $z_{a}$,
$S(\tilde{s}(-z_{1},\dots,-(z_{a}-\kappa),\dots,-z_{n}),s(z_{1},\dots,z_{a}-\kappa,\dots,z_{n}))=S(\tilde{s}(-z),s(z)).$
See also Corollary 7.10.
3 Pochhammer polynomials
For $\kappa\in{\mathbb{K}}$ and $m$ a positive integer, define Pochhammer polynomial $(t;\kappa)_{m}\in{\mathbb{K}}[t]$ by the formula $(t;\kappa)_{m}=\prod_{i=1}^{m}(t-(i-1)\kappa)$}. [^1] See https://en.wikipedia.org/wiki/Falling_and_rising_factorials. We have
$(t-\kappa;\kappa)_{m}=(t;\kappa)_{m}\frac{t-\kappa m}{t},\qquad(t+\kappa;\kappa)_{m}=(t;\kappa)_{m}\frac{t+\kappa}{t-(m-1)\kappa}.$
$(t+z;\kappa)_{m}=\sum_{i=0}^{m}{m\choose i}(t;\kappa)_{i}(z;\kappa)_{m-i},$
$(t;\kappa)_{i}(t;\kappa)_{j}=\sum_{l=0}^{\min(i,j)}\binom{i}{l}\binom{j}{l}l!\kappa^{l}(t;\kappa)_{i+j-l},$ (3.1)
$(t;\kappa)_{m}=\sum_{l=0}^{m}s_{1}(m,l)\kappa^{m-l}t^{l},$
$t^{m}=\sum_{l=0}^{m}s_{2}(m,l)\kappa^{m-l}(t;\kappa)_{l},$ (3.2)
where the integers $s_{1}(m,l)$ and $s_{2}(m,l)$ are Stirling numbers of the first and second kind, respectively. Notice that
$s_{1}(m,m)=s_{2}(m,m)=1.$ (3.3)
We also have $(t;\kappa)_{p}=t^{p}-\kappa^{p-1}t$,
$(t+z;\kappa)_{p}=(t+z)^{p}-\kappa^{p-1}(t+z)=t^{p}-\kappa^{p-1}t+z^{p}-\kappa^{p-1}z=(t;\kappa)_{p}+(z;\kappa)_{p}.$
We call a polynomial $f(t)\in{\mathbb{F}}_{p}[t]$ a quasi-constant if $f(t-\kappa)=f(t)$. The quasi-constants are polynomials in $t^{p}-\kappa^{p-1}t$. A Pochhammer polynomial $(t;\kappa)_{m}$ is a quasi-constant if $m$ is divisible by $p$. Then $(t;\kappa)_{pa}=\big(t^{p}-\kappa^{p-1}t\big)^{a}$.
Let $A$ be a $\mathbb{K}$}-algebra, for example, $A={\mathbb{K}}[z_{1},\dots,z_{n}]$. The polynomials $\{(t;\kappa)_{m}\mid m\geqslant 0\}$ form an $A$-basis of the ring of polynomials $A[t]$.
4 $\boldsymbol{p}$-hypergeometric solutions for $\boldsymbol{\kappa\in{\mathbb{F}}_{p}^{\times}\subset{\mathbb{K}}}$
4.1 Solutions in $\boldsymbol{\operatorname{Sing}L^{\otimes n}[n-2]}$
In this paper, we study solutions of the qKZ equations with values in $V:=\operatorname{Sing}L^{\otimes n}[n-2]$. The space $L^{\otimes n}[n-2]$ has a basis $v^{(i)}=v_{1}\otimes\dots\otimes v_{2}\otimes\dots\otimes v_{1}$, $i=1,\dots,n$, where the only $v_{2}$ stays at the $i$-th place. In this basis, the subspace $V$ consists of all vectors with the sum of coordinates equal to zero. We identify $L^{\otimes n}[n-2]$ with ${\mathbb{K}}^{n}$ and the subspace $V$ with the vector space
$\{x\in{\mathbb{K}}^{n}\mid x_{1}+\dots+x_{n}=0\}.$ (4.1)
4.2 Master polynomial and weight functions
For $\kappa\in{\mathbb{F}}_{p}^{\times}$, let $0<k<p$ be the positive integer such that
$\kappa k\equiv-1\pmod{p}$. (4.2)
Define master polynomial $\Phi(t,z;\kappa)=\prod_{a=1}^{n}(t-z_{a};\kappa)_{k}$, where $k$ is defined in (4.2). For $1\leqslant a\leqslant n$, define the weight functions
$\eta_{a}(t,z)=\frac{1}{t-z_{a}}\prod_{j=1}^{a-1}\frac{t-z_{j}+1}{t-z_{j}},\qquad Q_{a}(t,z;\kappa)=\Phi(t,z;\kappa)\eta_{a}(t,z).$
Notice that the formula for the function $\eta_{a}(t,z)$ does not depend on $p$ and $\kappa$.
Lemma 4.1.
The function $Q_{a}(t,z;\kappa)$ is a product of $n$ Pochhammer polynomials:
$Q_{a}(t,z;\kappa)=\left(\prod_{j=1}^{a-1}(t-z_{j}-\kappa;\kappa)_{k}\right)(t-z_{a}-\kappa;\kappa)_{k-1}\left(\prod_{j=a+1}^{n}(t-z_{j};\kappa)_{k}\right).$
Define a vector of polynomials
$Q(t,z;\kappa)=(Q_{1}(t,z;\kappa),\dots,Q_{n}(t,z;\kappa))^{\intercal}=\sum_{i}Q^{i}(z;\kappa)(t;\kappa)_{i},$ (4.3)
where $M^{\intercal}$ denotes the transpose matrix of a matrix $M$ and $Q^{i}(z;\kappa)=\big(Q^{i}_{1}(z;\kappa),\dots,Q^{i}_{n}(z;\kappa)\big)^{\intercal}$ are vectors of polynomials in $z$.
Example 4.2.
For $n=2$,
$Q(t,z;\kappa)=\big((t_{1}-z_{1}-\kappa;\kappa)_{k-1}(t_{1}-z_{2};\kappa)_{k},(t_{1}-z_{1}-\kappa;\kappa)_{k}(t_{1}-z_{2}-\kappa;\kappa)_{k-1}\big)^{\intercal}.$
Theorem 4.3 ([5, Theorem 5.1]).
For any positive integer $\ell$, the vector $Q^{\ell p-1}(z;\kappa)$ of polynomials in $z$ is a solutions of the qKZ equations with step $\kappa$ and values in $V=\operatorname{Sing}L^{\otimes n}[n-2]$, in particular, $\sum_{a=1}^{n}Q^{\ell p-1}_{a}(z;\kappa)=0$, see (4.1).
Let $[x]$ denote the integer part of a real number $x$. The vector $Q^{\ell p-1}(z;\kappa)$ is zero if $\ell\not\in\big\{1,\dots,\big[\frac{n{k}}{p}\big]\big\}$ for degree reasons. The vectors of polynomials $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,\big[\frac{n{k}}{p}\big]$, are called the $p$-hypergeometric solutions of the qKZ equations with step $\kappa$ and values in $V$. Denote $d(\kappa)=\big[\frac{n{k}}{p}\big]$}.
Notice that if $\frac{n{k}}{p}<1$, then there are no $p$-hypergeometric solutions.
4.3 Step $\boldsymbol{-\kappa}$
The integer $k$ satisfies the inequalities $0<k<p$ and the congruence $\kappa k\equiv-1\pmod{p}$, see (4.2). Then the integer pkp-k satisfies the inequalities 0<pk<p0<p-k<p and the congruence $\kappa(p-k)\equiv-1\pmod{p}$. Hence $\Phi(t,z;-\kappa)=\prod_{a=1}^{n}(t-z_{a};-\kappa)_{p-k}$. Recall that
$Q_{a}(t,z;-\kappa)=\Phi(t,z;-\kappa)\eta_{a}(t,z),$
$Q(t,z;-\kappa)=(Q_{1}(t,z;-\kappa),\dots,Q_{n}(t,z;-\kappa))^{\intercal}=\sum_{i}Q^{i}(z;-\kappa)(t;-\kappa)_{i}.$
Corollary 4.4.
The vectors
$Q^{\ell p-1}(z;-\kappa),\qquad\ell=1,\dots,\left[\frac{n(p-k)}{p}\right],$
are solutions of the qKZ equations with step $\kappa$ and values in $V$.
If $p$ does not divide $n$, then the total number of $p$-hypergeometric solutions of the qKZ equations with values in $V$ and steps $\kappa$ and $\kappa$ equals
$\left[\frac{n{k}}{p}\right]+\left[\frac{n(p-{k})}{p}\right]=n-1=\dim V.$
4.4 $\boldsymbol{p}$-hypergeometric solutions of KZ equations
In this subsection, we remind the construction in [6] of polynomial solutions modulo $p$ of the differential KZ equations with values in $V=\operatorname{Sing}L^{\otimes n}[n-2]$.
Let $0<k<p$ be the positive integers such that $\kappa k\equiv-1\pmod{p}$. Define master polynomial $\bar{\Phi}(t,z,\kappa)=\prod_{a=1}^{n}(t-z_{a})^{k}$}. For $1\leqslant a\leqslant n$, define the weight functions
$\bar{w}_{a}(t,z)=\frac{1}{t-z_{a}},\qquad\bar{Q}_{a}(t,z;\kappa)=\bar{\Phi}(t,z;\kappa)\bar{w}_{a}(t,z).$
Then $\bar{Q}_{a}(t,z;\kappa)$ is a polynomial in $t,z$. Define a vector of polynomials in $t$, $z$,
$\bar{Q}(t,z;\kappa)=\big(\bar{Q}_{1}(t,z;\kappa),\dots,\bar{Q}_{n}(t,z;\kappa)\big)^{\intercal}=\sum_{i}\bar{Q}^{i}(z;\kappa)t^{i},$
where $\bar{Q}^{i}(z;\kappa)=\big(\bar{Q}^{i}_{1}(z;\kappa),\dots,\bar{Q}^{i}_{n}(z;\kappa)\big)^{\intercal}$ are vectors of polynomials in $z$.
Theorem 4.5 ([6]).
For any positive integer $\ell$, the vector $\bar{Q}^{\ell p-1}(z;\kappa)$ is a solution of the differential KZ equations with parameter $\kappa$ and values in $V=\operatorname{Sing}L^{\otimes n}[n-2]$.
The vector $\bar{Q}^{\ell p-1}(z;\kappa)$ is zero if $\ell\not\in\big\{1,\dots,\big[\frac{n{k}}{p}\big]\big\}$ for degree reasons. The vectors
$\bar{Q}^{\ell p-1}(z,\kappa),\qquad\ell=1,\dots,\left[\frac{n{k}}{p}\right],$
are called the $p$-hypergeometric solutions of the differential KZ equations with parameter $\kappa$ and values in $V$.
Notice that $\bar{Q}_{a}(t,z;\kappa)$ are homogeneous polynomials in variables $t$, $z$ of degree nk1nk-1, and $\bar{Q}^{\ell p-1}(z;\kappa)$} are vectors of homogeneous polynomials in $z$ of degree $nk-\ell p$.
4.5 Top-degree part of $\boldsymbol{p}$-hypergeometric solutions
It turns out that the top-degree part of a $p$-hypergeometric solution $Q^{\ell p-1}(z;\kappa)$ of the qKZ equations is the $p$-hypergeometric solution $\bar{Q}^{\ell p-1}(z;\kappa)$ of the KZ equations.
We start with an abstract lemma. Let $t,z_{1},\dots,z_{n},\alpha$ be variables. Given a homogeneous polynomial $P(t,z_{1},\dots,z_{n},\alpha)$ of degree $d$ in the variables $t$, $z$, $\alpha$ and an integer $e$, $0\leqslant e\leqslant d$, we construct two polynomials $\bar{P}_{e}(z_{1},\dots,z_{n})$ and $P_{e}(z_{1},\dots,z_{n},0)$ as follows.
On the one hand, we have
$P(t,z_{1},\dots,z_{n},\alpha)=\sum_{d_{0}+\dots+d_{n+1}=d}a_{d_{0},\dots,d_{n+1}}t^{d_{0}}z_{1}^{d_{1}}\cdots z_{n}^{d_{n}}\alpha^{d_{n+1}}.$ (4.4)
Then
$P(t,z_{1},\dots,z_{n},0)=\sum_{d_{0}+\dots+d_{n}=d}a_{d_{0},\dots,d_{n},0}t^{d_{0}}z_{1}^{d_{1}}\cdots z_{n}^{d_{n}}.$ (4.5)
Denote
$\bar{P}_{e}(z_{1},\dots,z_{n})=\sum_{e+d_{1}+\dots+d_{n}=d}a_{e,d_{1},\dots,d_{n},0}z_{1}^{d_{1}}\cdots z_{n}^{d_{n}},$
the coefficient of $t^{e}$ in (4.5).
On the other hand, applying formula (3.2) to each $t^{d_{0}}$ at (4.4), we rewrite this sum as
$P(t,z_{1},\dots,z_{n},\alpha)=\sum_{d_{0}+\dots+d_{n+1}=d}b_{d_{0},\dots,d_{n+1}}(t;\alpha)_{d_{0}}z_{1}^{d_{1}}\cdots z_{n}^{d_{n}}\alpha^{d_{n+1}}.$ (4.6)
Denote
$P_{e}(z_{1},\dots,z_{n},\alpha)=\sum_{e+d_{1}+\dots+d_{n+1}=d}b_{e,d_{1},\dots,d_{n+1}}z_{1}^{d_{1}}\cdots z_{n}^{d_{n}}\alpha^{d_{n+1}},$
the coefficient of $(t;\alpha)_{e}$ in (4.6). Then
$P_{e}(z_{1},\dots,z_{n},0)=\sum_{e+d_{1}+\dots+d_{n}=d}b_{e,d_{1},\dots,d_{n},0}z_{1}^{d_{1}}\cdots z_{n}^{d_{n}}.$
Lemma 4.6.
We have $P_{e}(z_{1},\dots,z_{n},0)=\bar{P}_{e}(z_{1},\dots,z_{n})$.
Proof.
By formula (3.3), we have $a_{e,d_{1},\dots,d_{n},0}=b_{e,d_{1},\dots,d_{n},0}$ for any $e,d_{1},\dots,d_{n}$. This implies the lemma. ∎
Corollary 4.7 ([5]).
Let $Q^{\ell p-1}(z;\kappa)$ be a $p$-hypergeometric solution of the qKZ equations from Theorem 4.3. Then $Q^{\ell p-1}(z;\kappa)=\bar{Q}^{\ell p-1}(z;\kappa)+\cdots$, where $\bar{Q}^{\ell p-1}(z;\kappa)$ is the corresponding $p$-hypergeometric solution of the KZ equations from Theorem 4.5, and the dots denote the terms of degree less than $nk-\ell p=\deg\bar{Q}^{\ell p-1}(z;\kappa)$.
Proof.
The corollary is proved by application of Lemma 4.6 to the polynomial $Q(t,z;\kappa)$ and integer $e=\ell p-1$. Then it is easy to see that $\bar{P}_{\ell p-1}(z_{1},\dots,z_{n})=\bar{Q}^{\ell p-1}(z;\kappa)$ and $P_{\ell p-1}(z_{1},\dots,z_{n},0)$ is the top-degree part of $Q^{\ell p-1}(z;\kappa)$. ∎
5 Linear independence
5.1 Lexicographical order
Define length-lexicographical order on monomials $v=z_{1}^{d_{1}}\cdots z_{n}^{d_{n}}$: $v<w$ if degree(v)<(v)<degree(w)(w) or if degree(v)=(v)=degree(w)(w) and $v$ is smaller than $w$ in lexicographical order.
For a polynomial $f(z)=\sum_{d_{1},\dots,d_{n}}a_{d_{1},\dots,d_{n}}z_{1}^{d_{1}}\cdots z_{n}^{d_{n}}$ denote by $\text{L}f(z)$ the nonzero summand $a_{d_{1},\dots,d_{n}}z_{1}^{d_{1}}\cdots z_{n}^{d_{n}}$ with the lexicographically largest monomial $z_{1}^{d_{1}}\cdots z_{n}^{d_{n}}$. We call $\text{L}f(z)$ the leading term of $f(z)$.
5.2 Leading terms
For $\ell\in\{1,\dots,d(\kappa)\}$, let $r(\ell)$ be the unique non-negative integer such that
$r(\ell)k\leqslant nk-\ell p<(r(\ell)+1)k.$
We have $r(1)>r(2)>\cdots$. Denote
$ga=(n-r(\ell))k-\ell p,\qquad u_{\ell}=\frac{(-1)^{nk-\ell p}}{k}\binom{k}{a}(0,\dots,0,k-a,k,\dots,k)^{\intercal}\in{\mathbb{F}}_{p}^{n},$
where kak-a is the $r(\ell)+1$-st coordinate. Notice that the integers $k$ and kak-a are not divisible by $p$, and $u_{\ell}\in{\mathbb{F}}_{p}^{n}$ is a singular vector since the sum of its coordinates equals zero.
Lemma 5.1 ([7, Lemma 3.1]).
Assume that $p$ does not divide $n$ and $\ell\in\{1,\dots,d(\kappa)\}$. Then the leading term of $\bar{Q}^{\ell p-1}(z;\kappa)$ is given by the formula
$\operatorname{L}\bar{Q}^{\ell p-1}(z;\kappa)=(z_{1}\cdots z_{r(\ell)})^{k}z_{r(\ell)+1}^{a}u_{\ell}.$
Lemma 5.2.
Assume that $p$ does not divide $n$ and $\ell\in\{1,\dots,d(\kappa)\}$. Let $Q^{\ell p-1}(z;\kappa)$ be the corresponding $p$-hypergeometric solution of the qKZ equations and $\bar{Q}^{\ell p-1}(z;\kappa)$ the corresponding $p$-hypergeometric solution of the KZ equations. Then their leading terms are equal,
$\operatorname{L}Q^{\ell p-1}(z;\kappa)=\operatorname{L}\bar{Q}^{\ell p-1}(z;\kappa).$
Proof.
The lemma follows from Corollary 4.7. ∎
Corollary 5.3.
If $p$ does not divide $n$ and $\ell\in\{1,\dots,d(\kappa)\}$, then the leading term of $Q^{\ell p-1}(z;\kappa)$ is given by the formula
$\operatorname{L}Q^{\ell p-1}(z;\kappa)=(z_{1}\cdots z_{r(\ell)})^{k}z_{r(\ell)+1}^{a}u_{\ell}.$ (5.1)
Example 5.4.
Let $n=3$ and $d(\kappa)=\big[\frac{3k}{p}\big]=1$. Then $\frac{p}{3}<k<\frac{2p}{3}$. In this case there is exactly one $p$-hypergeometric solution $Q^{p-1}(z;\kappa)$. If $\frac{p}{2}<k<\frac{2p}{3}$, then the leading term of $Q^{p-1}(z;\kappa)$ is
$z_{1}^{k}z_{2}^{2k-p}\frac{(-1)^{3k-p}}{k}\binom{k}{2k-p}(0,p-k,k)^{\intercal}.$
If $\frac{p}{3}<k<\frac{p}{2}$, then the leading term of $Q^{p-1}(z;\kappa)$ is
$z_{1}^{3k-p}\frac{(-1)^{k}}{k}\binom{k}{3k-p}(p-2k,k,k)^{\intercal}.$
Consider the collections of the $n$-vectors $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$, as an $(n\times d(\kappa))$-matrix. For $I=\{1\leqslant i_{1}<\dots<i_{d(\kappa)}\leqslant n\}$, denote by $M_{I}(z,\kappa)$ the $(d(\kappa)\times d(\kappa))$-minor of that matrix located at the rows with indices in $I$.
Lemma 5.5.
If $p$ does not divide $n$ and $I=\{r(d(\kappa))<\dots<r(1)\}$, then the minor $M_{I}(z;\kappa)$ is a nonzero polynomial.
Proof.
The lemma is a corollary of formula (5.1). ∎
Theorem 5.6.
If $p$ does not divide $n$, then the $p$-hypergeometric solutions $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$, of the qKZ equations are linearly independent over the field ${\mathbb{K}}(z)$.
Proof.
The statement follows from formula (5.1) and the fact that the vectors $u_{\ell}$ are linearly independent over $\mathbb{K}$}. ∎
6 Orthogonality relations
6.1 Statement
Recall that $d(-\kappa)=\big[\frac{n(p-k)}{p}\big]$.
Theorem 6.1.
Let $p>n$ and $0<d(\kappa)<n-1$. Then for any $\ell\in\{1,\dots,d(\kappa)\}$ and $m\in\{1,\dots,d(-\kappa)\}$, we have
$S\big(Q^{mp-1}(-z;-\kappa),Q^{\ell p-1}(z;\kappa)\big)=\sum_{a=1}^{n}Q^{mp-1}_{a}(-z;-\kappa)Q^{\ell p-1}_{a}(z;\kappa)=0,$ (6.1)
where $S$ is the Shapovalov form.
The theorem is proved in Sections 6.2 and 6.3.
Remark.
It is easy to see that the Shapovalov form on $V$ is nondegenerate if $p$ does not divide $n$. Indeed, the vectors $e_{1}=(1,-1,0,\dots,0)$, $e_{2}=(0,1,-1,0,\dots,0)$, …, $e_{n-1}=(0,\dots,0,1,-1)$ form a basis of $V$, and the determinant of the Shapovalov form in this basis equals $n$.
Remark.
Formula (6.1) and Corollary 4.7 imply the orthogonality relations for the $p$-hypergeometric solutions of the KZ equations,
$S\big(\bar{Q}^{mp-1}(-z;-\kappa),\bar{Q}^{\ell p-1}(z;\kappa)\big)=\sum_{a=1}^{n}\bar{Q}^{mp-1}_{a}(-z;-\kappa)\bar{Q}^{\ell p-1}_{a}(z;\kappa).=0.$ (6.2)
Two different proofs of formula (6.2) are given in [8, Theorem 3.11] and [8, Appendix A].
6.2 Special restrictions
Let $I\subset\{1,\dots,n\}$ be a nonempty subset, $I=\{1\leqslant i_{1}<i_{2}<\dots<i_{a}\leqslant n\}$. Denote $S_{I}$ the system of equations
$z_{i_{b}}=((b-1)k-1)\kappa,\qquad b=1,\dots,a.$ (6.3)
For a polynomial $f(z)$, define $f(z)_{S_{I}}$ to be the polynomial $f(z)$ in which the variables $(z_{i})_{i\in I}$ are replaced by multiples of $\kappa$ according to formulas (6.3).
Lemma 6.2.
Let $Q(t,z;\kappa)$ be the vector of polynomials defined by (4.3). Let $I=\{1\leqslant i_{1}<i_{2}<\dots<i_{a}\leqslant n\}\subset\{1,\dots,n\}$ be a nonempty subset. Then
$Q(t,z;\kappa)_{S_{I}}=(t;\kappa)_{ak-1}(P_{1}(t,z),\dots,P_{n}(t,z))^{\intercal},$
where $P_{1}(t,z),\dots,P_{n}(t,z)$ are suitable polynomials of degree (na)k(n-a)k.
Proof.
The proof is straightforward. ∎
Corollary 6.3.
Let $Q^{\ell p-1}(z;\kappa)$ be a $p$-hypergeometric solution and $\ell p<ak$. Then
$Q^{\ell p-1}(z;\kappa)_{S_{I}}=0.$ (6.4)
Proof.
For $j=1,\dots,n$, the $j$-th coordinate of $Q(t,z;\kappa)_{S_{I}}$ equals $(t;\kappa)_{ak-1}P_{j}(t,z)$. Using (3.1), we rewrite this as $\sum_{i\geqslant ak-1}c_{i}(z)(t;\kappa)_{i}$ for suitable $c_{i}(z)$. We observe that $(t;\kappa)_{\ell p-1}$ does not enter this sum. This proves the corollary. ∎
6.3 Proof of Theorem 6.1
Denote
$G_{\ell,m}(z;\kappa)=S\big(Q^{mp-1}(-z;-\kappa),Q^{\ell p-1}(z;\kappa)\big).$
Then
$G_{\ell,m}(-z;\kappa)={}$ $S\big(Q^{mp-1}(z;-\kappa),Q^{\ell p-1}(-z;\kappa)\big)$
$={}$ $S\big(Q^{\ell p-1}(-z;\kappa),Q^{mp-1}(z;-\kappa)\big)=G_{m,\ell}(z;-\kappa).$
Hence, $G_{\ell,m}(z;\kappa)=0$ for all $\ell$, $m$ if and only if $G_{m,\ell}(z;-\kappa)=0$.
By Lemma 2.4, the function $G_{\ell,m}(z;\kappa)$ is a polynomial in ${\mathbb{F}}_{p}[z_{1}^{p}-z_{1},\dots,z_{n}^{p}-z_{n}]$. Denote $h(x)=x^{p}-x$.
Lemma 6.4.
Given $\ell$ and $m$, we have
$G_{\ell,m}(z;\kappa)=c_{0}+\sum_{b=1}^{n-\ell-m}\sum_{1\leqslant j_{1}<\dots<j_{b}\leqslant n}c_{j_{1},\dots,j_{b}}h(z_{j_{1}})\cdots h(z_{j_{b}}),\qquad c_{0},c_{j_{1},\dots,j_{b}}\in{\mathbb{F}}_{p}.$
Proof.
On the one hand, we have
$\deg Q^{\ell p-1}(z;\kappa)=kn-\ell p\qquad\text{and}\qquad\deg Q^{mp-1}(-z;-\kappa)=(p-k)n-mp.$
Hence $\deg G_{\ell,m}(z;\kappa)\leqslant(n-\ell-m)p$.
On the other hand, for any $j=1,\dots,n$, we have $\deg_{z_{j}}G_{\ell,m}(z;\kappa)\leqslant p$ since $\deg_{z_{j}}Q(t,z;\kappa)=k$ and $\deg_{z_{j}}Q(t,-z;-\kappa)=p-k$. These two remarks prove the lemma. ∎
The $p$-hypergeometric solution $Q^{\ell p-1}(z,\kappa)$ is associated with the integer $0<k<p$ such that $\kappa k\equiv-1\pmod{p}$}, while the $p$-hypergeometric solution $Q^{mp-1}(z,-\kappa)$ is associated with the integer 0<pk<p0<p-k<p such that $\kappa(p-k)\equiv-1\pmod{p}$. Given $m$, $\ell$, we say that $k$ is a good parameter if $\ell(p-k)<mk$. We say that pkp-k is a good parameter if $mk<\ell(p-k)$. Notice that $mk\neq\ell(p-k)$. Otherwise $p$ must divide $\ell+m$ which is impossible since $\ell+m\leqslant n-1<p$.
Having the two integers $k$, pkp-k and two solutions $Q^{\ell p-1}(z;\kappa)$ and $Q^{mp-1}(z;-\kappa)$, we may and will assume that $k$ denotes the good parameter.
Lemma 6.5.
Given $Q^{\ell p-1}(z;\kappa)$ and $Q^{mp-1}(z;-\kappa)$, assume that $k$ is a good parameter. Then $G_{\ell,m}(z;\kappa)=0$.
Proof.
Let $\{1,\dots,n\}=I\cup J$ be a partition where $|I|=\ell+m$ and $|J|=n-\ell-m$. We may apply formula (6.4) to $Q^{\ell p-1}(z;\kappa)$ with $a=\ell+m$ since $k$ is a good parameter and hence $\ell p<(\ell+m)k$. Hence we have $G_{\ell,m}(z;\kappa)_{I}=0$, where
$G_{\ell,m}(z;\kappa)_{I}=c_{0}+\sum_{b=1}^{n-\ell-m}\sum_{\{j_{1}<\dots<j_{b}\}\subset J}c_{j_{1},\dots,j_{b}}h(z_{j_{1}})\dots h(z_{j_{b}}).$ (6.5)
Hence the right-hand side polynomial at (6.5) is the zero polynomial for any subset II with $|I|=\ell+m$. Therefore, $G_{\ell,m}(z;\kappa)$ is the zero polynomial. ∎
Lemma 6.5 implies Theorem 6.1.
7 Invariant subbundles
7.1 Subbundle of qKZ connection
Assume that $p>n$ and $\kappa\in{\mathbb{F}}_{p}^{\times}$. Consider the discrete qKZ connection on the direct product $V\times{\mathbb{K}}^{n}\to{\mathbb{K}}^{n}$}, where $V=\{x\in{\mathbb{K}}^{n}\mid x_{1}+\dots+x_{n}=0\}$ and the qKZ operators are defined by formula (2.2),
$K_{a}(z;\kappa)=R^{(a,a-1)}(z_{a}-z_{a-1}-\kappa)\cdots R^{(a,1)}(z_{a}-z_{1}-\kappa)$
[…]
The connection has singularities at the points of ${\mathbb{K}}^{n}$ where the qKZ operators have poles or are degenerate.
The $R$-matrix $R(u)=\frac{u-P}{u-1}$ has a pole if $u=1$ and is degenerate if u=1u=-1. Denote by $H_{i,j,m}$ the affine hyperplane in ${\mathbb{K}}^{n}$ defined by the equation $z_{i}-z_{j}-m=0$ where $1\leqslant i<j\leqslant n$, $m\in{\mathbb{F}}_{p}$. Let $\bar{\mathcal{A}}^{\circ}$ be the arrangement in ${\mathbb{K}}^{n}$ of all hyperplanes $H_{i,j,m}$. Let $\bar{\mathcal{A}}={\mathbb{K}}^{n}-\bar{\mathcal{A}}^{\circ}$ denote its complement. The qKZ operators are well-defined over $\bar{\mathcal{A}}$ and are nondegenerate.
Assume that $0<d(\kappa)<n-1$. Then the $p$-hypergeometric solutions $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$, define flat sections of the qKZ connection which we call the $p$-hypergeometric sections.
Recall the minors $M_{I}(z;\kappa)$ defined for any $I=\{1\leqslant i_{1}<\dots<i_{d(\kappa)}\leqslant n\}$ in Section 5.2 with the help of these $p$-hypergeometric sections. Denote by $\mathcal{A}(\kappa)$ the Zariski open subset of $\bar{\mathcal{A}}$ consisting of points $b\in\bar{\mathcal{A}}$ such that at least one of the minors $M_{I}(z;\kappa)$ is nonzero at $b$.
For any point $b\in\mathcal{A}(\kappa)$, the vectors $Q^{\ell p-1}(b,\kappa)$, $\ell=1,\dots,d(\kappa)$, are linearly independent and span a $d(\kappa)$-dimensional $\mathbb{K}$}-vector subspace $\mathcal{S}(b,\kappa)$ of $V$. These subspaces $\mathcal{S}(b,\kappa)$, $b\in\mathcal{A}(\kappa)$, form a vector subbundle $\mathcal{S}(\kappa)\to\mathcal{A}(\kappa)$ of the trivial bundle $V\times\mathcal{A}(\kappa)\to\mathcal{A}(\kappa)$.
Remark.
Notice that the minors $M_{I}(z;\kappa)$ are polynomials in $z$ with coefficients in ${\mathbb{F}}_{p}$ and are independent of the field $\mathbb{K}$}. Notice also that the base $\mathcal{A}(\kappa)$ is invariant with respect to the affine translations, $(z_{1},\dots,z_{n})\mapsto(z_{1},\dots,z_{a}-\kappa,\dots,z_{n})$, $a=1,\dots,n$.
The subbundle $\mathcal{S}(\kappa)\to\mathcal{A}(\kappa)$ is invariant under the qKZ connection, and the $p$-hypergeometric sections form a flat basis of the space of its sections.
We also consider the quotient bundle $\mathcal{Q}(\kappa)\to\mathcal{A}(\kappa)$ with fibers $V/\mathcal{S}(b,\kappa)$. The qKZ connection on $V\times\mathcal{A}(\kappa)\to\mathcal{A}(\kappa)$ induces a discrete connection on $\mathcal{Q}(\kappa)\to\mathcal{A}(\kappa)$ which we also call the qKZ connection. Notice that the rank of $\mathcal{Q}(\kappa)\to\mathcal{A}(\kappa)$ equals
$\dim V-\operatorname{rank}\mathcal{S}(\kappa)=n-1-\left[\frac{n{k}}{p}\right]=\left[\frac{n{(p-k)}}{p}\right]=d(-\kappa).$
If $d(\kappa)=0$, we define $\mathcal{A}(\kappa)=\bar{\mathcal{A}}$. In this case, we define $\mathcal{S}(\kappa)\to\mathcal{A}(\kappa)$ to be the rank 0 subbundle of $V\times\mathcal{A}(\kappa)$ and also define $\mathcal{Q}(\kappa)\to\mathcal{A}(\kappa)$ to be $V\times\mathcal{A}(\kappa)$.
7.2 Subbundle of dual qKZ connection
Assume that $p>n$ and $\kappa\in{\mathbb{F}}_{p}^{\times}$. Consider the dual discrete qKZ connection on $V\times{\mathbb{K}}^{n}\to{\mathbb{K}}^{n}$ defined by formulas (2.11) and (2.7). The dual qKZ operators $K_{a}(-z_{1},\dots,-z_{n};-\kappa)$, $a=1,\dots,n$, are well-defined over $\bar{\mathcal{A}}$ and are nondegenerate.
Assume that $0<d(-\kappa)<n-1$. This assumption is equivalent to the assumption $0<d(\kappa)<n-1$. Consider the $p$-hypergeometric solutions $Q^{mp-1}(z;-\kappa)$, $m=1,\dots,d(-\kappa)$, of the qKZ equations step $\kappa$. Then the $V$-valued polynomials $Q^{mp-1}(-z;-\kappa)$, $m=1,\dots,d(-\kappa)$, define flat sections of the dual qKZ connection which we also call the $p$-hypergeometric sections.
Recall the minors $M_{I}(z;-\kappa)$ defined for any $I=\{1\leqslant i_{1}<\dots<i_{d(-\kappa)}\leqslant n\}$ in Section 5.2 with the help of the $p$-hypergeometric solutions $Q^{mp-1}(z;-\kappa)$, $m=1,\dots,d(-\kappa)$. Denote by $\mathcal{B}(\kappa)$ the Zariski open subset of $\bar{\mathcal{A}}$ consisting of points $b\in\bar{\mathcal{A}}$ such that at least one of the polynomials $M_{I}(-z;-\kappa)$ is nonzero at $b$.
For any point $b\in\mathcal{B}(\kappa)$, the vectors $Q^{mp-1}(-b,-\kappa)$, $m=1,\dots,d(-\kappa)$, are linearly independent and span a $d(-\kappa)$-dimensional $\mathbb{K}$}-vector subspace $\mathcal{S}^{*}(b,-\kappa)$ of space $V$. These subspaces $\mathcal{S}^{*}(b,-\kappa)$, $b\in\mathcal{B}(\kappa)$, form a vector subbundle $\mathcal{S}^{*}(\kappa)\to\mathcal{B}(\kappa)$ of the trivial bundle $V\times\mathcal{B}(\kappa)\to\mathcal{B}(\kappa)$}.
The subbundle $\mathcal{S}^{*}(\kappa)\to\mathcal{B}(\kappa)$ is invariant with respect to the dual qKZ connection, and the $p$-hypergeometric sections $Q^{mp-1}(-z;-\kappa)$, $m=1,\dots,d(-\kappa)$, form a flat basis of the space of its sections.
Consider the restriction of the bundles $\mathcal{S}^{*}(\kappa)\to\mathcal{B}(\kappa)$ and $\mathcal{Q}(\kappa)\to\mathcal{A}(\kappa)$ to $\mathcal{A}(\kappa)\cap\mathcal{B}(\kappa)$. For any $b\in\mathcal{A}(\kappa)\cap\mathcal{B}(\kappa)$, the Shapovalov form defines a nondegenerate pairing
$S\colon\ \mathcal{S}^{*}(b,-\kappa)\otimes V/\mathcal{S}(b,\kappa)\to{\mathbb{K}}$
of the fibers of these bundles, by Theorem 6.1. The discrete connections on $\mathcal{S}^{*}(\kappa)\to\mathcal{B}(\kappa)$ and $\mathcal{Q}(\kappa)\to\mathcal{A}(\kappa)$} are dual with respect to the Shapovalov form, that is, for any $u\in\mathcal{S}^{*}(b,-\kappa)$, $v\in V/\mathcal{S}(b,\kappa)$, and $a=1,\dots,n$, we have $S(u,v)=S(K_{a}(-b;-\kappa)u,K_{a}(b;\kappa)v)$.
Define $p$-quasi-hypergeometric sections $T^{\ell}(z;\kappa)$, $\ell=1,\dots,d(-\kappa)$, of the bundle $\mathcal{Q}(\kappa)\to\mathcal{A}(\kappa)$ over $\mathcal{A}(\kappa)\cap\mathcal{B}(\kappa)$ by the formulas
$S\big(Q^{mp-1}(-b,-\kappa),T^{\ell}(b,\kappa)\big)=\delta_{\ell,m},\qquad m=1,\dots,d(-\kappa).$
Lemma 7.1.
Assume that $p>n$ and $0<d(-\kappa)\leqslant n-1$. Then $p$-quasi-hypergeometric sections $T^{\ell}(z;\kappa)$}, $\ell=1,\dots,d(-\kappa)$, of the quotient bundle $\mathcal{Q}(\kappa)\to\mathcal{A}(\kappa)\cap\mathcal{B}(\kappa)$ form a flat basis of the space of sections of that bundle.
Proof.
The proof is straightforward. ∎
If $d(-\kappa)=0$, we define $\mathcal{B}(\kappa)=\bar{\mathcal{A}}$. We also define $\mathcal{S}^{*}(\kappa)\to\mathcal{B}(\kappa)$ to be the rank 0 subbundle of $V\times\mathcal{B}(\kappa)$.
Example 7.2.
Let $n=2$. Then $V$ is of dimension 1. For $p=5$, $k=\kappa=3$, we have $d(3)=1$, and the qKZ connection has a flat basis given by the $p$-hypergeometric solution
$Q^{4}(z_{1},z_{2})=(-2z_{1}+2z_{2}+2,2z_{1}-2z_{2}-2).$
For $p=5$. $k=\kappa=2$, we have $d(2)=0$, and the qKZ connection has a flat basis given by the $p$-quasi-hypergeometric solution
$T^{1}(z_{1},z_{2},2)=\left(\frac{3}{2z_{1}-2z_{2}+2},\frac{3}{-2z_{1}+2z_{2}-2}\right).$
7.3 Reduced $\boldsymbol{p}$-curvature operators
Let $\hat{C}_{a}(z;\kappa)$, $a=1,\dots,n$, be the reduced $p$-curvature operators of the qKZ discrete connection on $V\times{\mathbb{K}}^{n}\to{\mathbb{K}}^{n}$, see (2.5).
Theorem 7.3.
If $p>n$, $\kappa\in{\mathbb{F}}_{p}^{\times}$, and $0<d(\kappa)<n-1$, then the span of the $p$-hypergeometric sections $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$, lies in the kernel of $\hat{C}_{a}(z,\kappa)$ and contains the image of $\hat{C}_{a}(z;\kappa)$ for every $a=1,\dots,n$.
If $d(\kappa)=p-1$ or 0, then all reduced $p$-curvature operators equal zero.
Corollary 7.4.
We have $\hat{C}_{a}(z;\kappa)\hat{C}_{b}(z;\kappa)=0$, $a,b=1,\dots,n$.
Proof of Theorem 7.3..
The span lies in the kernel of $\hat{C}_{a}(z;\kappa)$ by formula (2.4).
The operator $\hat{C}_{a}(z;\kappa)$ annihilates the span, hence $\hat{C}_{a}(z;\kappa)$ induces a well-defined operator on the fibers of the quotient bundle $\mathcal{Q}(\kappa)\to\mathcal{A}(\kappa)$. This induced operator is the $a$-th reduced $p$-curvature operator of the qKZ connection on the quotient bundle. The quotient bundle has a flat basis of $p$-quasi-hypergeometric sections over the Zariski open subset $\mathcal{A}(\kappa)\cap\mathcal{B}(\kappa)$. Hence all reduced $p$-curvature operators of the qKZ connection on the quotient bundle are zero. Therefore the image of $\hat{C}_{a}(z;\kappa)$ is contained in the span.
If $d(\kappa)=p-1$, then $\mathcal{S}(\kappa)\to\mathcal{A}(\kappa)$ coincides with $V\times\mathcal{A}(\kappa)\to\mathcal{A}(\kappa)$, and all reduced $p$-curvature operators are zero by formula (2.4).
If $d(\kappa)=0$, then $p$-quasi-hypergeometric sections form a flat basis of the space of sections of $V\times\mathcal{A}(\kappa)\to\mathcal{A}(\kappa)$}, and again all reduced $p$-curvature operators are zero by formula (2.4). ∎
Lemma 7.5.
If $p>n$, $\kappa\in{\mathbb{F}}_{p}^{\times}$, and $0<d(\kappa)<n-1$, then every reduced $p$-curvature operator $\tilde{C}_{a}(z;\kappa)$} is nonzero.
Proof.
Consider a normalized $p$-curvature operator $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$ of the associated differential KZ equations. In a basis of $V$, the entries of the matrix of the operator $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$ are homogeneous polynomials in $z$ of degree (n2)p(n-2)p. By Corollary 2.2, in the same basis, the entries of the matrix of the operator $\tilde{C}_{a}(z,\kappa)$ are polynomials in $z$ of degree (n2)p(n-2)p whose top-degree parts equal the corresponding entries of the matrix of the operator $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$.
It is proved in [8, Theorem 1.13] that if $p>n$, $\kappa\in{\mathbb{F}}_{p}^{\times}$, then every reduced $p$-curvature operator $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$ is a (nonzero) operator of rank 1. Hence every reduced $p$-curvature operator $\tilde{C}_{a}(z,\kappa)$ is a nonzero operator. ∎
Example 7.6.
For $n=3$, we have $\dim V=2$. Let $p>3$ and $d(\kappa)=1$. Then $d(-\kappa)=1$. For $a=1,2,3$, the kernel of the reduced $p$-curvature operator $\hat{C}_{a}(z,\kappa)$ is generated by $Q^{p-1}(z,\kappa)$ and the image of $\hat{C}_{a}(z,\kappa)$ is generated by $Q^{p-1}(z,\kappa)$. Such an operator is determined uniquely up to multiplication by a scalar rational function in $z$.
For an operator $F\colon V\to V$, denote by $F^{*}\colon V\to V$ the operator dual to $F$ under the Shapovalov form, $S(F^{*}x,y)=S(x,Fy)$.
Lemma 7.7.
We have
$\hat{C}_{a}(z,-\kappa)=-\hat{C}_{a}(-z;\kappa)^{*}.$ (7.1)
Proof.
We have $S(x,y)=S(C_{a}(-z;-\kappa)x,C_{a}(z;\kappa)y)$ by formulas (2.10) and (2.7). Hence $C_{a}(-z;-\kappa)=\big(C_{a}(z;\kappa)^{-1}\big)^{*}$. We also have $(C_{a}(z;\kappa)-1)^{2}=0$ by Corollary 7.4. Then
$C_{a}(z;\kappa)^{-1}=(1+(C_{a}(z;\kappa)-1))^{-1}=1-(C_{a}(z;\kappa)-1)$
and $C_{a}(-z;-\kappa)-1=1-C_{a}(z;\kappa)^{*}$. ∎
Corollary 7.8.
The normalized $p$-curvature operators satisfy the equation
$\tilde{C}_{a}(z;-\kappa)=(-1)^{n}\tilde{C}_{a}(-z;\kappa)^{*}.$ (7.2)
Proof.
Formula (7.2) follows from equation (7.1) and the following formulas:
$\tilde{C}_{a}(z;\kappa)=(C_{a}(z;\kappa)-1)\prod_{j\neq i}(z_{i}-z_{j},\kappa)_{p},$
$\tilde{C}_{a}(-z;-\kappa)=(C_{a}(-z;-\kappa)-1)\prod_{j\neq i}(-z_{i}+z_{j},-\kappa)_{p}.$
7.4 All solutions of qKZ equations for $\boldsymbol{\kappa\in{\mathbb{F}}_{p}^{\times}}$
Theorem 7.9.
Let $p>n$, $\kappa\in{\mathbb{F}}_{p}^{\times}$, and $0<d(\kappa)<n-1$. Let $f(z)$ be a $V$-valued rational function in $z$ which is a solution of the qKZ equations with step $\kappa$. Then $f(z)$ is a linear combination of the $p$-hypergeometric solutions $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$, with coefficients which are rational functions in $z_{i}^{p}-z_{i}$, $i=1,\dots,n$.
Recall that if $d(\kappa)=n-1$, then the qKZ connection has a basis of flat sections given by the $p$-hypergeometric sections by Theorem 5.6, and if $d(\kappa)=0$, then the qKZ connection has a basis of flat sections given by the $p$-quasi-hypergeometric solutions, by Lemma 7.1.
Proof.
For $a=1,\dots,n$, consider the normalized $p$-curvature operators $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$ and $\tilde{C}_{a}(z,\kappa)$. Both of these operators are polynomials in $z$, and the polynomial $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$ is the top-degree part of the polynomial $\tilde{C}_{a}(z,\kappa)$. The polynomial $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$ is nonzero by [8, Theorem 1.13] and hence $\tilde{C}_{a}(z,\kappa)$ is a nonzero operator.
It was proved in [8, Theorem 1.8] that if $p>n$, $\kappa\in{\mathbb{F}}_{p}^{\times}$, and $0<d(\kappa)<n-1$, then all solutions of the KZ equations are linear combinations of the $p$-hypergeometric solutions. Hence the intersection of kernels of the operators $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$, $a=1,\dots,n$, is of dimension $d(\kappa)$ for generic $z$, and the span of images of the operators $\tilde{C}_{a}^{\operatorname{KZ}}(z,\kappa)$, $a=1,\dots,n$, is of dimension $n-1-d(\kappa)$} for generic $z$. Therefore, the span of images of the operators $\tilde{C}_{a}(z,\kappa)$, $a=1,\dots,n$, has dimension at least $n-1-d(\kappa)$ for generic $z$. This implies that the span of values of flat sections of the qKZ connection is of dimension not larger than $d(\kappa)$ for generic $z$. But we have $d(\kappa)$ flat linear independent $p$-hypergeometric sections $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,d(\kappa)$. Hence any flat section of the qKZ connection is a linear combination of the $p$-hypergeometric sections with 1-periodic coefficients. ∎
Corollary 7.10.
Let $p>n$, $\kappa\in{\mathbb{F}}_{p}^{\times}$, and $0<d(\kappa)<n-1$. Let $f(z)$ and $g(z)$ be $V$-valued rational functions in $z$ where $f(z)$ is a solution the qKZ equations with step $\kappa$ and $g(z)$ is a solution the qKZ equations with step $\kappa$. Then
$S(g(-z),f(z))=0.$ (7.3)
Formula (7.3) follows from Theorems 7.9 and 6.1.
7.5 qKZ connection with $\boldsymbol{\kappa\in{\mathbb{K}}\setminus{\mathbb{F}}_{p}}$
Lemma 7.11.
Let $p>n$ and $\kappa\in{\mathbb{K}}\setminus{\mathbb{F}}_{p}$ . Then all the normalized $p$-curvature operators $\tilde{C}_{a}(z;\kappa)$}, $a=1,\dots,n$, are nondegenerate for generic $z$.
Proof.
Formula (3.19) in [8] describes the spectrum of the $p$-curvature operators $C_{a}^{\operatorname{KZ}}(z;\kappa)$ of differential KZ equations. The formula shows that all $p$-curvature operators $C_{a}^{\operatorname{KZ}}(z;\kappa)$ are nondegenerate for generic $z$. In a basis of $V$, the matrices of $\tilde{C}_{a}^{\operatorname{KZ}}(z;\kappa)$ are homogeneous polynomials in $z$ of degree (n2)p(n-2)p. Hence their determinants are nonzero homogeneous polynomials in $z$. By Corollary 2.2, the determinants of the normalized $p$-curvature operators $\tilde{C}_{a}(z;\kappa)$ are nonzero polynomials. The lemma follows. ∎
Corollary 7.12.
For $p>n$ and $\kappa\in{\mathbb{K}}\setminus{\mathbb{F}}_{p}$, there does not exist a nonzero rational $V$-valued function $f(z)$ which is a flat section of the qKZ connection with parameter $\kappa$.
Proof.
If $f(z)$ is a flat section, then it lies in the kernel of every normalized $p$-curvature operator $\tilde{C}_{a}(z;\kappa)$. That contradicts to Lemma 7.11. ∎
Acknowledgements
E.M. was supported in part by Simons Foundation grant 709444. A.V. was supported in part by Simons Foundation grant TSM-00012774. The authors thank P. Etingof, R. Rimanyi, and V. Vologodsky for useful discussions.